a wall is supported by a brace 10 feet long. if one end of the brace is placed 6 feet from the base of the wall, how man feet up the wall does the brace reach?
Answer by nerdybill(6959)

a wall is supported by a brace 10 feet long. if one end of the brace is placed 6 feet from the base of the wall, how man feet up the wall does the brace reach?
.
Since the wall and the floor forms a right angle (90 degrees), we can apply the Pythagorean theorem (where the brace is the hypotenuse).
.
Let x = distance (feet) the brace reaches up the wall
then
x^2 + 6^2 = 10^2
x^2 + 36 = 100
x^2 = 64
x = 8 feet